OpenStudy (anonymous):

The sum of two nonnegative numbers is 36. Find these numbers if the first plus the square of the second is a maximum.

OpenStudy (anonymous):

Set up an equation. What are you given?

OpenStudy (anonymous):

that's all the information i was given and i have no idea what to do! :(

OpenStudy (asnaseer):

Hero: it should be:\[x+y=36\]

hero (hero):

x + y = 36 (duh) x + y^2 = maximum (Wondering how to figure out the maximum)

OpenStudy (anonymous):

can someone please help me figure out the answer

OpenStudy (turingtest):

is this calculus?

OpenStudy (anonymous):

yes it is

hero (hero):

Yeah, this definitely would involve calculus. Taking a derivative would help.

hero (hero):

m = x + (36 - x)^2

OpenStudy (turingtest):

express the second equation in terms of x, then set the derivative equal to zero

OpenStudy (anonymous):

x+y=36 x+y^2 = maximum it can be inferred by ituition that y=36 and x= 0 gives the maximum, although for a rigorous proof replace x in terms of y, get y^2-y+36 upward opening parabola, no local maxima exists, hence we take the largest value in the domain, which is y=36. hence the solution is x=0 and y=36

OpenStudy (anonymous):

setting the derivative as zero will give minima and not maxima @TuringTest

OpenStudy (anonymous):

would it be 6 instead of 36 since a number has to be squared and 6 squared is 36?

OpenStudy (turingtest):

setting the derivative equal to zero will give any extrema, max or min @Stom

OpenStudy (anonymous):

its 36 and not 6 , as the question says that sum of those 2 nimbers is 36, @turingtest ya,

OpenStudy (turingtest):

...so in this case, the maximum occurred at one of the endpoints when looking for an absolute max or min, you must check all extrema (i.e. when the derivative equals zero) as well as the endpoints of the possible intervals. In this case x=0 y=36 and x=36 y=0 are the endpoints